what is the slope intercepy form of the linear equation with a graph that passes through (3,-5) and is perpendicular to the graph of the followin equation?
2x-3y=9

Answer by ewatrrr(24785)


Hi
Using the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept.
Line Perpendicular to 2x-3y=9 and passing thru Pt(3,-5)
2x - 3y = 9 Or y = (-2/3)x -3 slope m = (-2/3)
New Line
y = (3/2)x + b
-5 = (3/2)*-2 + b
-2 = b
y = (3/2)x - 2 (Blue Line)